Jennifer takes 4 hours to do a job.john takes 6 hours to do the same job.working together,how many hours will it take them to do the job?
Answer by JulietG(1812)

Jennifer finishes the work at a rate of 1/4 per hour
John finishes the work at a rate of 1/6 per hour
After one hour, 1/4 + 1/6 is finished, or 3/12 + 2/12 -> 5/12
.
If they do the job at the rate of 5/12 per hour, how long will it take to complete 12/12?
12/5 = 2.4 hours
.
Let's check.
2.4 hours for Jennifer at 1/4 per hour = 2.4 * .25 = .6 of the job
2.4 hours for John at 1/6 per hour = 2.4 * 1/6 = .4 of the job
.6 + .4 = 1.0
Success!